. A monopolist has the following weekly total revenue and total cost function (R) = 30Q- Q^2 and (C )= Q^3 -15Q^2+10Q+100, respectively in dollars.


a) Find the level of output that maximizes the profit?


b) Find the maximum weekly profit?


c) Find the point elasticity of demand at equilibrium level of output?

1
Expert's answer
2022-01-16T17:50:47-0500

a) The level of output that maximizes the profit is:

MR = MC,

MR = R'(Q) = 30 - 2Q,

"MC = C'(Q) = 3Q^2 - 30Q +10,"

"3Q^2 - 30Q +10 = 30 - 2Q,"

"3Q^2 - 28Q - 20 = 0,"

Q = (28 + 32)/6 = 10 units.


b) The maximum weekly profit is:

"TP = R - C = (30\u00d710 - 10^2) - (10^3 - 15\u00d710^2 + 10\u00d710 +100) = 500."


c) The point elasticity of demand at equilibrium level of output is:

P = 30 - Q = 30 - 10 = 20,

Q = 30 - P, then:

Ed = -1×20/10 = -2, so the demand is elastic.